Two cars are moving. The first car has thrice the mass of the second car but only 1/3 as the 

kinetic energy. When both cars increase their speed by 9 m/s, they then have the same kinetic 

energy. 

Calculate the original speeds of the first car.


1
Expert's answer
2021-04-29T10:43:24-0400
"M=3m\\0.5Mv^2=0.5\frac{1}{3}mV^2\\3v^2=\frac{1}{3}V^2\\V=3v"

"0.5M(v+9)^2=0.5m(V+9)^2\\3(v+9)^2=(3v+9)^2\\v=5.2\frac{m}{s}\\V=15.6\frac{m}{s}"